OpenStudy (anonymous):

6^2 ÷ 6 × 4 + 6 A) 12 B) 30 C) 18 D) 144 Use order of operations, I keep getting 7.5. PEMDAS

OpenStudy (anonymous):

In the Calculator, I get 30.

OpenStudy (anonymous):

The equation is (6^2)/6*4+6. So 36/6*4+6 = 6*4+6 = 30 6 to the power of 2 has the higest priority here

OpenStudy (anonymous):

I did 6^2 first, which is E, exponents, then the next is M, which is Multiply, and i get 24, then next comes D, which is divide, and I divide 36/24 and get 1.5, then A for Add it to 6.

OpenStudy (anonymous):

Im using PEMDAS.

OpenStudy (anonymous):

its 30. btw the calculator follows pedmas too

OpenStudy (anonymous):

I can't use a calculator on the test. :P

OpenStudy (anonymous):

winston if there are multiply's and divide's in an equation do the operations from left to right for only multiply and divide part. Otherwise you'll find a wrong answer. First 36/6, than 6*4. okay?

OpenStudy (anonymous):

you are doing 6^2 ÷ (6 x 4) + 6 which is wrong it is not dividing by 24, it is dividing by 6 and then multiplying by 4. what you are doing is equivalent to 6^2 ÷ 6 ÷ 4 + 6

OpenStudy (anonymous):

Euler is right